40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 16261     Entry time: Tue Jul 27 23:04:37 2021
Author: Anchal 
Type: Update 
Category: LSC 
Subject: 40 meter party 

[ian, anchal, paco]

After our second attempt of locking PRFPMI tonight, we tried to resotre XARM and YARM locks to IR by clicking on IFO_CONFIGURE>Restore XARM (POX) and IFO_CONFIGURE>Restore YARM (POY) but the arms did not lock. The green lasers were locked to the arms at maximum power, so the relative alignments of each cavity was ok. We were also able to lock PRMI using IFO_CONFIGURE>Restore PRMI carrier.

This was very weird to us. We were pretty sure that the aligment is correct, so we decided to cehck the POX POY signal chain. There was essentially no signal coming at POX11 and there was a -100 offset on it. We could see some PDH signal on POY11 but not enough to catch the locks.

We tried running IFO_CONFIGURE>LSC OFFSETS to cancel out any dark current DC offsets. The changes made by the script are shown in attachment 1.

We went to check the tables and found no light visible on beam finder cards on POX11 or POY11. We found that ITMX was stuck on one of the coils. We unstuck it using the shaking method. The OPLEVs on ITMX after this could not be switched on as the OPLEV servo were railing to limits. But when we ran Restore XARM (POX) again, they started working fine. Something is done by this script that we are not aware of.

We're stopping here. We still can not lock any of the single arms.


Wed Jul 28 11:19:00 2021 Update:

[gautam, paco]

Gautam found that the restoring of POX/POY failed to restore the whitening filter gains in POX11 / POY11. These are meant to be restored to 30 dB and 18 dB for POX11 and POY11 respectively but were set to 0 dB in detriment of any POX/POY triggering/locking. The reason these are lowered is to avoid saturating the speakers during lock acquisition. Yesterday, burt-restore didn't work because we restored the c1lscepics.snap but said gains are actually in c1lscaux.snap. After manually restoring the POX11 and POY11 whitening filter gains, gautam ran the LSCOffsets script. The XARM and YARM were able to quickly lock after we restored these settings.

The root of our issue may be that we didn't run the CARM & DARM watch script (which can be accessed from the ALS/Watch Scripts in medm). Gautam added a line on the Transition_IR_ALS.py script to run the watch script instead.

Attachment 1: Screenshot_2021-07-27_22-19-58.png  33 kB  Uploaded Wed Jul 28 00:16:04 2021  | Hide | Hide all
Screenshot_2021-07-27_22-19-58.png
ELOG V3.1.3-